the function f(r)=πr^2 determines the area of a circle in cm2, A, in terms of the radius of the circle in cm, r.

a. Write a function g that determines the radius of a circle in cm, r, in terms of the area of the cirlce in cm2, A.
g(A)=
b. Use function notation to answer the following questions:
-What is the area (in cm2) of a circle whose radius is 2.21 cm?
-What is the radius (in cm) of a circle whose area is 1798 cm2?
-How much greater is the area of a circle whose radius is 15.3 cm than the area of a circle whose radius is 15.2 cm?

Answer by Boreal(15235)

area=pi*r^2 so sqrt (area/pi)=r
g(r)=sqrt)area/pi
A=pi*2.21^2=15.34 cm^2
r=sqrt(1798/pi)=23.92 cm
it is the ratio of (15.3/15.2)^2=1.013 times greater
15.3^2*pi=735.42
15.2^*pi=725.83
That ratio is 1.013